OpenStudy (anonymous):

Solve the quadratic equation: 2x2 + 11x − 6 = 0

OpenStudy (anonymous):

2x^2 + 11x -6 = 0 (2x - 1) (x + 6) A^2 + Bx + C = 0 Find two numbers that have a product of C and add/Subtract the numbers to get B. So start out by finding two factors that equal 6. In this case, By plugging in 6 and 1 into the factor, it would multiply out as 2x^2 +12x - x -6 12x - x = 11x so the two factors work.